7. An 80.0 kg student running at 3.5 m/s grabs onto a rope that it hanging vertically.
a. How much kinetic energy does the student have prior to grabbing the rope?
b. How high will the student swing?
1
Expert's answer
2020-08-10T12:32:19-0400

a. "E_k={{mv^2}\over2}={{80\sdot{3.5^2}}\over2}=490" J

b. "E_k=E_p\implies" "{{mv^2}\over2}=mgh\implies" "h={{v^2}\over2g}={{{3.5^2}}\over2\sdot9.81}=0.624" m = 62.4 cm